# Butterfly in Spacetime: Inherent Instabilities in Stable Black Holes

### v2

@@ Abstract
This paper numerically studies if the stability of a stable black hole is robustness against the small perturbation on geometry near its event horizon. As a toy model, it encodes the such perturbation into deformations of Regge-Wheeler potential. It considers three different types of local deformations-the negative static bump potential, the stochastic potential and bump potential modulated by time function in low frequency limit. Our numerical results show that infinitesimal local deformations on Regge-Wheeler potential near the horizon can overturn stability of an arbitrarily strongly stable black hole, implying that late-time behavior of a stable black hole is extremely sensitive to geometry near horizon. Specially, certain deformations that stabilize systems in flat backgrounds can destabilize otherwise stable black holes. It also shows that horizon-induced redshift transforms near-horizon quantum fluctuations into classical-scale stochastic deformations capable of triggering instability, implying that even an isolated black hole cannot keep stable in extended timescales.
@@ I Introduction
In the field of black hole (BH) physics, the analysis of BH stability stands as a fundamental and critical subject, as it reveals whether BHs can maintain their existence in the universe over extended periods. The quasi-normal modes (QNMs), meanwhile, serve as one of the essential physical quantities for exploring and understanding the stability characteristics of BHsChandrasekhar and Detweiler (1975); Nollert and Price (1999); Kokkotas and Schmidt (1999); Berti et al. (2009); Konoplya and Zhidenko (2011). Unlike normal modes in close system, the QNMs are a series of discrete complex modes. The sign of the imaginary parts indicates whether the spacetime is stable or not. In general, QNMs indicating unstable spacetime describe a system where phase transition happens, such as spontaneous scalarization of compact objectSilva et al. (2018); Herdeiro et al. (2018); Doneva et al. (2024); Xu et al. (2024); Xiong and Li (2024), while QNMs describing stable spacetime play an important role in ringdown signal of gravitational waves in the last period of two astrophysical compact object mergeCreighton (1999); Dreyer et al. (2004); Berti et al. (2006); Gossan et al. (2012); Berti et al. (2018); Wang et al. (2024).
@@
Recently, an interesting feature of QNMs, even the fundamental modes, has been reported that they are sensitive to small alterations of background, which called spectrum instabilityNollert (1996); Cheung et al. (2022); Cardoso et al. (2024); Qian et al. (2025). For example, when the effective potential is altered by adding a small positive “bump” away from the location of photon ring, the quasi-normal modes (QNMs) of a Schwarzschild BH are completely alteredCheung et al. (2022). The sensitivity of QNMs seems to imply that the QNMs are not appropriate observables in ringdown stage. However, the ringdown signal appear to be robust by the spectral instabilityDaghigh et al. (2020); Berti et al. (2022); Yang et al. (2024); Ianniccari et al. (2024); Rosato et al. (2024). During the early ringdown stage, the signal in altered background exhibits no substantial divergence from the original signal. In the late ringdown stage, while a discernible discrepancy arises between the two, the magnitude of this difference is directly proportional to the intensity of the bump and will exponentially decays to zero. Notably, both signals undergo rapid decay to zero, maintaining consistency in their asymptotic behavior. The ringdown signal is nonsensitive against the small alterations of effective potential in these literatures.
@@
When addressing the topic of BH stability, the focus is usually on whether a probe field exhibits exponential growth in a fixed spacetime background. Numerous studies have demonstrated that Schwarzschild BH remains stable under perturbations from scalar fields, Dirac fields, and spin-2 fieldsRegge and Wheeler (1957); Chandrasekhar (1998). Due to the complex environment and unavoidable classical/quantum fluctuations of real BHs, no models in these studies can fully characterize any real BH in the universe. However, these investigations are still regarded as highly valuable. A key reason is the prevailing belief that the conclusion about BH stability is robust against small alterations of the spacetime metric-that is, sufficiently small alteration on the metric would not disrupt the stability of an initially stable BH. In other words, if a theoretical BH described by metric $g_{\mu\nu}$ is stable, then we believe the real BH described by metric $\tilde{g}_{\mu\nu}$ will also be stable if the difference between $\tilde{g}_{\mu\nu}$ and $g_{\mu\nu}$ is “small” enough. However, our results in this work suggest that this “taken-for-granted” assumption may not be valid. Using the Schwarzschild BH as a concrete example, we consider if its stability is sensitive to the modification of geometry near horizon. As a toy model, we encode such modification into the deformation of Regge-Wheeler potential. We analyze the long-time behavior of probe scalar field when a small negative or zero-mean random local effective potential is added sufficiently close to the horizon, extending the analysis beyond the positive bump scenario investigated inCheung et al. (2022); Destounis and Duque (2024). In the time domain, we numerically show that the spacetime becomes unstable when the deformation is sufficiently close to the horizon. Particularly, certain perturbed potentials that stabilize systems in flat backgrounds can destabilize otherwise stable black holes. We conclude that, at least under static spherical symmetry, the stability of a BH does not exhibit robustness. Two nearly identical BHs may exhibit completely different stability properties at late times. It will also show that horizon-induced redshift transforms near-horizon quantum fluctuations into classical-scale stochastic local deformations on the metric, which can trigger instability of the BH and cause the BH gradually to lose energy over time. In this paper, we use natural units $\hbar=c=G=1$.
II Effective potential and its deformations
@@
$\text{d}s^{2}=-h(r)\text{d}t^{2}+\frac{\text{d}r^{2}}{h(r)}+r^{2}\text{d}\Omega^{2},~{}~{}~{}h(r)=1-r_{h}/r\,,$ (1)
where $r_{h}$ is the radius of event horizon, $\text{d}\Omega^{2}$ stands for line element of unit sphere. In time domain, we consider equation of motion with respect to the s-wave scalar perturbation $\psi(t,r_{*})$ in the Schwarzschild spacetime
@@
We now add a deformation term into the effective potential to capture possible environmental effects presenting around the BH system. Detailed discussion on how the environments change Eq. (2) will much involved and we here simply assume that they are all encoded into a deformation on effective potential
@@
In following discussions, we will specify $r_{h}=1$. We present our numerical results for solving Eq. (2) in time domain with initial data corresponding to Gaussian wave packet
$\psi(0,r_{*})=\exp\left(\frac{r_{*}^{2}}{20}\right)\,,\quad\frac{\partial\psi}{\partial t}(0,r_{*})=0\,.$ (4)
For spatial grid $r_{*}$, we adopt forth-order central difference method for spatial derivative and integrate the Eq. (2) using forth-order Rugge-Kutta method for $t$. Here we will consider three different types of potential $V_{p}(r_{*}-a,t)$.
@@ III Static negative bump
The first one is the static “bump” potential $V_{p}(x,t)=V_{p}(x)$, where $V_{p}(x)$ has certain sign and compact support. When $V_{p}(x)$ is positive, it has been well known that such deformation will lead to “spectrum instability”. However, this kind of “instability” is not the instability of BH itself, i.e., the system will still settle down into a static BH without scalar hair. Since the bumps are added by hand, there is not proper justification for a positive or negative sign. Particularly, near the event horizon, the quantum effects of gravity and matter may lead to exotic matter and may raise negative “bump”. The negative “bump” can even appear if we impose dominate energy condition (see one example in appendix). The negative "bump" could resemble a tachyonic field Cardoso et al. (2009); Baibhav et al. (2023) or be inherent in various modified gravity theories, such as the Einstein-Gauss-Bonnet theory Cao et al. (2024). If the bump is close to a stable BH, the situation become complicated. We expect a competition with the absorption at the horizon. For strong enough negative bump, this feature might give rise to spacetime instabilitiesCardoso et al. (2009); Baibhav et al. (2023); Cao et al. (2024). Naively, one may expect that the BH will be still stable if the original BH is strongly enough but the negative “bump” is small enough. Surprisingly, our numerical results show that, if the infinitesimal negative “bump” is close to event horizon enough, it can turnover arbitrarily stable BH.
@@
We numerically investigate how this negative bump change the long-time behavior of scalar perturbation when it is placed close to even horizon.
Figure 1: Left panel: Time evolution of $\psi$ at $x=a$ for different values of $\epsilon$ and $a$. Here $t_{0}$ is defined as the first time when $\psi(a,t)$ is at its local maximum value. Right panel: Relationship between $\epsilon$ and $|a_{c}|$. The error bars are determined by the bounds of bisection. The red line is given by equation $\ln|a_{c}|=-\ln|\overline{V}_{p}|-\ln\epsilon$ (not by fitting).
The left panel of Fig. 1. shows the results as we move perturbed potential towards event horizon by choosing $\epsilon=10^{-2}$ and $4\times 10^{-3}$. It shows that for small value of $a$ where the deformations are confined near the photon ring, $\psi$ will decay to approach zero in the last stage, confirming that the BH is stable. However, for the case $\epsilon=10^{-2}$ we found that there exist a critical distance $a_{c}\approx-15$ of deformation to trigger the BH instability. When $a<a_{c}$, the perturbation field $\psi$ will experience exponential growth at late time. This exponential amplification signals an onset of instability, indicating that the perturbation field will drive the BH towards an unstable state during the later stage of its evolution.
@@
In the second form we suppose $V_{p}(x,t)$ to be local stochastic function with zero mean. This in some sense is closer to real astrophysical environments compared with a fixed-sign perturbation. Many astrophysical BHs are surrounded by dynamical evolutional matter. The matter supplies non-static stochastic deformations on the geometry of BH. As a toy model, the effects of random matter then could be mimicked by random non-static modification on the Regge-Wheeler potential. More importantly, this kind of modification seemingly is inevitably even for an isolated BH without any matter surrounding it, since it can come from the local quantum fluctuations. Based on equivalent principle, the wavelength $\lambda_{p}$ and period $t_{p}$ of local quantum fluctuation measured by local inertial observers are Planck scales $\lambda_{p}=t_{p}\sim\mathcal{O}(l_{\rm P})$. These quantum fluctuations will lead that the geometry of spacetime varies in the local proper time scale $t_{p}$ and local proper distance scale $\lambda_{p}$. Usually, such extreme ultraviolet variaions of metric have no dramatic effect on classical physics. But the appearance of horizon makes thing different due to the strong redshift.
To illustrate that, we consider the quantum fluctuation happening in the near-event-horizon region $r=r_{h}+\delta r$ (measured in coordinate gauge (1)). The line element in terms of tortoise coordinate in this region then will be $\text{d}s^{2}=h(r)(-\text{d}t^{2}+\text{d}r_{*}^{2})\approx h^{\prime}(r_{h})\delta r(-\text{d}t^{2}+\text{d}r_{*}^{2})$ such that one finds $\lambda_{p}=\sqrt{h^{\prime}(r_{h})\delta r}\lambda=\sqrt{h^{\prime}(r_{h})\delta r}T$. Here $\lambda$ and $T$ denote the wavelength and period in tortoise coordinate. The picture of classical horizon would fail if one is close to it near Planck scale. Thus, the proper distance between $r=r_{h}+\delta r$ and $r=r_{h}$ can at most as same as Planck scale. This means $\delta r/\sqrt{h(r)}\approx\delta r/\sqrt{h^{\prime}(r_{h})\delta r}\sim\mathcal{O}(\lambda_{p})$. One thus find that $T=\lambda\sim 1/h^{\prime}(r_{h})\sim\mathcal{O}(r_{h})$. Therefore, from the perspective of tortoise coordinate, near horizon local quantum fluctuations would induce metric or Regge-Wheeler potential to varies in infrared scale. It needs to emphasize again that, in the case of absent horizon or in the region far way from horizon, the local quantum fluctuations will have no essential contributions on classical physics in infrared scale.
@@
where $W(x)$ is random function localizing around $x=0$ and $U(t)$ is a temporal random function. The function $W$ and $U$ satisfy
@@
Figure 2: Left panel: Time evolution of $\psi$ at $x=a$ for different values of $\epsilon$ and $a$. Right panel: Relationship between $\epsilon$ and $|a_{c}|$. The red line is $\ln|a_{c}|=-2\ln\epsilon+0.02$, which is obtained by fitting the most left 7 points.

### v3

@@ Abstract
This paper numerically studies if the stability of a stable black hole is robust against the small perturbation on geometry near its event horizon. As a toy model, it encodes the such perturbation into deformations of Regge-Wheeler potential. It considers three different types of local deformations-the negative static bump potential, the stochastic potential and bump potential modulated by time function in low frequency limit. Our numerical results show that infinitesimal local deformations on Regge-Wheeler potential near the horizon can overturn stability of a stable black hole, implying that late-time behavior of a stable black hole is extremely sensitive to geometry near horizon. Specially, certain deformations that stabilize systems in flat backgrounds can destabilize otherwise stable black holes. It also shows that horizon-induced redshift transforms near-horizon quantum fluctuations into classical-scale stochastic deformations capable of triggering instability, implying that even an isolated black hole cannot keep stable in extended timescales.
@@ I Introduction
In the field of black hole (BH) physics, the analysis of BH stability stands as a fundamental and critical subject, as it reveals whether BHs can maintain their existence in the universe over extended periods. The stability of a BH addresses the question: if it is subjected to a small perturbation, will the perturbations decay with time, allowing the BH to return to its original state, or will they grow uncontrollably and destroy the configuration altogether? The stability problem has deep implications for the mathematical structure of general relativity, cosmic censorship, and the no-hair conjecture.Chandrasekhar (1983). The quasi-normal modes (QNMs), meanwhile, serve as one of the essential physical quantities for exploring and understanding the stability characteristics of BHsChandrasekhar and Detweiler (1975); Nollert and Price (1999); Kokkotas and Schmidt (1999); Berti et al. (2009); Konoplya and Zhidenko (2011). Unlike normal modes in closed system, the QNMs are a series of discrete complex modes. The sign of the imaginary parts indicates whether the spacetime is stable or not. In general, QNMs indicating unstable spacetime describe a system where phase transition happens, such as spontaneous scalarization of compact objectSilva et al. (2018); Herdeiro et al. (2018); Doneva et al. (2024); Xu et al. (2024); Xiong and Li (2024), while QNMs describing stable spacetime play an important role in ringdown signal of gravitational waves in the last period of two astrophysical compact object mergerCreighton (1999); Dreyer et al. (2004); Berti et al. (2006); Gossan et al. (2012); Berti et al. (2018); Wang et al. (2024).
@@
Recently, an interesting feature of QNMs, even the fundamental modes, has been reported that they are sensitive to small alterations of background, which is called spectrum instabilityNollert (1996); Cheung et al. (2022); Cardoso et al. (2024); Qian et al. (2025). For example, when the effective potential is altered by adding a small positive “bump” away from the location of photon ring, the quasi-normal modes (QNMs) of a Schwarzschild BH are completely alteredCheung et al. (2022).In addition, the spiral formula of spectrum instability can be semi-analytically derived in double rectangular barriers using perturbation theory (See the Supplemental material in Motohashi (2025)) The sensitivity of QNMs seems to imply that the QNMs are not appropriate observables in ringdown stage. However, the ringdown signal appears to be robust by the spectral instability Daghigh et al. (2020); Kyutoku et al. (2023); Berti et al. (2022); Yang et al. (2024); Ianniccari et al. (2024); Rosato et al. (2024). Specifically, it was pointed out that the spectrum instability can hardly affect the greybody factor and on the ringdown signal. Further, the phase shift of the greybody factor could carry more physical information when studying the environmental effects on QNMs and the ringdown signal Kyutoku et al. (2023).
During the early ringdown stage, the signal in altered background exhibits no substantial divergence from the original signal. In the late ringdown stage, while a discernible discrepancy arises between the two, the magnitude of this difference is directly proportional to the intensity of the bump and will exponentially decays to zero. Notably, both signals undergo rapid decay to zero, maintaining consistency in their asymptotic behavior. The ringdown signal is nonsensitive against the small positive alterations of effective potential. From a more macroscopic perspective, the stability of BHs remains unchallenged by this minor alteration on effective potential.
@@
When addressing the topic of BH stability, the focus is linear stability of BHs, which usually focuses on whether a probe field exhibits exponential growth in a fixed spacetime background. Linear stability analysis of BHs involves solving wave-like equations for perturbations of the metric or matter fields Chandrasekhar (1983). Numerous studies have demonstrated that Schwarzschild BH remains stable under perturbations from scalar fields, Dirac fields, and spin-2 fieldsRegge and Wheeler (1957); Chandrasekhar (1998). Due to the complex environment and unavoidable classical/quantum fluctuations of real BHs, no models in these studies can fully characterize any real BH in the universe. However, these investigations are still regarded as highly valuable. A key reason is the prevailing belief that the conclusion about BH stability is robust against small alterations of the spacetime metric-that is, sufficiently small alteration on the metric would not disrupt the stability of an initially stable BH. In other words, if a theoretical BH described by metric $g_{\mu\nu}$ is stable, then we believe the real BH described by metric $\tilde{g}_{\mu\nu}$ will also be stable if the difference between $\tilde{g}_{\mu\nu}$ and $g_{\mu\nu}$ is “small” enough. This naive idea is seemingly supported by some models in studying “spectrum instability” Destounis and Duque (2024), since the perturbed potential in those models do not change instability of BHs in time domain.
However, our results in this work suggest that this “taken-for-granted” assumption may not be valid. Using the Schwarzschild BH as a concrete example, we consider if its stability is sensitive to the modification of geometry near horizon. As a toy model, we encode such modification into the alteration of Regge-Wheeler potential. We analyze the long-time behavior of probe scalar field when a small negative or zero-mean random local effective potential is added sufficiently close to the horizon, extending the analysis beyond the positive bump scenario investigated inCheung et al. (2022). In the time domain, we numerically show that the spacetime becomes unstable when the alteration is sufficiently close to the horizon. Particularly, certain perturbed potentials that stabilize systems in flat backgrounds can destabilize otherwise stable BHs. We conclude that, at least under static spherical symmetry, BH stability does not exhibit robustness. Two nearly identical BHs may exhibit completely different stability properties at late time. It will also show that horizon-induced redshift transforms near-horizon quantum fluctuations into classical-scale stochastic local deformations on the metric, which can trigger instability of the BH and cause it gradually to lose energy over time. In this paper, we use natural units $\hbar=c=G=1$.
II Effective potential and its alterations
@@
$\text{d}s^{2}=-h\text{d}t^{2}+\frac{\text{d}r^{2}}{h}+r^{2}\text{d}\Omega^{2}\,$, (1)
$h=1-\frac{r_{h}}{r}\,$, (2)
where $r_{h}$ is the radius of event horizon, $\text{d}\Omega^{2}$ stands for line element of unit sphere. In time domain, we then consider equation of motion with respect to the s-wave scalar perturbation $\Psi=\psi(t,r_{*})/r$ as a test fieldin the Schwarzschild spacetime
@@
We now add an alteration term into the effective potential to capture possible environmental effects presenting around the BH system. Detailed discussion on how the environments change Eq. (3) would be much involved and we here simply assume that they are all encoded into an alteration on effective potential
@@
In the following discussions, we will specify $r_{h}=1$. We present our numerical results for solving Eq. (3) in time domain with initial data corresponding to Gaussian wave packet
$\psi(r_{*},0)=\exp\left(-\frac{r_{*}^{2}}{20}\right)\,,\quad\frac{\partial\psi}{\partial t}(r_{*},0)=0\,.$ (5)
For spatial grid $r_{*}$, we adopt fourth-order central difference method for spatial derivative and integrate the Eq. (3) using forth-order Rugge-Kutta method for $t$. At the horizon ($r_{*}\rightarrow$-$\infty$) and the infinity ($r_{*}\rightarrow\infty$) we impose the no reflect boundary conditions. Here we will consider three different types of potential $V_{p}(r_{*}-a,t)$.
@@ III Static negative bump
The first one is the “bump” static potential $V_{p}(x,t)=V_{p}(x)$, where $V_{p}(x)$ has certain sign and compact support. When $V_{p}(x)$ is positive, it has been well known that such alteration will lead to “spectrum instability”. However, this kind of “instability” is not the instability of BH itself, i.e., the system will still settle down into a static BH without scalar hair. Since the bumps are added by hand, there is no proper justification for a positive or negative sign. Particularly, near the event horizon, the quantum effects of gravity and matter may lead to exotic matter and may raise negative “bump”. The negative “bump” can even appear if we impose dominant energy condition (see one example in appendix). The negative "bump" could resemble a tachyonic field or be inherent in various modified gravity theories, such as the Einstein-Gauss-Bonnet theory. This feature might give rise to spacetime instabilities Cardoso et al. (2009); Baibhav et al. (2023); Cao et al. (2024). If the bump is close to a stable BH, the situation becomes complicated. We expect a competition with the absorption at the horizon. Naively, one may expect that the BH will be still stable if the original BH is strongly enough but the negative “bump” is small enough. Surprisingly, our numerical results show that, if the infinitesimal negative “bump” is close to event horizon enough, it can overturn an arbitrarily stable BH.
@@
We numerically investigate how this negative bump changes the long-time behavior of scalar perturbation when it is placed close to even horizon.
Figure 1: Left panel: Time evolutions of $\psi$ at $x=a$ for different values of $\epsilon$ and $a$. Here $t_{0}$ is defined as the first time at which $\psi(a,t)$ is at its local maximum value. Right panel: Relationship between $\epsilon$ and $|a_{c}|$. The error bars are determined by the bounds of bisection. The red line is given by equation $\ln|a_{c}|=-\ln|\overline{V}_{p}|-\ln\epsilon$ (not by fitting).
The left panel of Fig. 1. shows the results as we move perturbed potential towards event horizon by choosing $\epsilon=10^{-2}$ and $4\times 10^{-3}$. It shows that for small value of $a$ where the alteration are confined near the photon ring, $\psi$ will decay to approach zero in the last stage, confirming that the Schwarzschild BH is stable. However, for the cases $\epsilon=10^{-2}$ we found that there exist a critical distance $a_{c}\approx-15$ of alteration to trigger the BH instability. When $a<a_{c}$, the perturbation field $\psi$ will experience exponential growth at late time. This exponential amplification signals an onset of instability, indicating that the perturbation field will drive the Schwarzschild BH towards an unstable state during the later stage of its evolution.
@@
The “negative bump” we discussed here could destabilize the flat spacetime independently. However, the existence of BH will absorb energy and stabilize the spacetime. Thus, roughly speaking, the instability can be understood as the competition between the destabilizing caused by perturbed potential $V_{p}$ and the stabilizing caused by BH. One may expect that as long as the “negative bump” is sufficiently small and sufficiently close to the BH, the stable BH should dominate the stability of spacetime configuration, leading that entire spacetime configuration should tend to be stable. However, our results present a counterintuitive conclusion: no matter how small the bump is, it can overturn the stability of the entire spacetime, turning an originally stable Schwarzschild BH into an unstable one. Particularly, in the appendix B, we consider the situation that the background is not a BH but a star, then we find that infinitesimal negative bump cannot trigger the instability if it is placed sufficiently close to the star. Thus, the instability discovered here not only is caused by negative bump but also involves intrinsic property of the black hole geometry.
@@
In the second form we suppose $V_{p}(x,t)$ to be local stochastic function with zero mean. This in some sense is closer to real astrophysical environments compared with a fixed-sign perturbation. Many astrophysical BHs are surrounded by dynamical evolutional matter. The matter supplies non-static stochastic deformations on the geometry of BH. As a toy model, the effects of random matter then could be mimicked by random non-static modification on the Regge-Wheeler potential. More importantly, this kind of modification seemingly is inevitably even for an isolated BH without any matter surrounding it, since it can come from the local quantum fluctuations. Based on the equivalent principle, the wavelength $\lambda_{p}$ and period $t_{p}$ of local quantum fluctuation measured by local inertial observers are Planck scale $\lambda_{p}=t_{p}\sim\mathcal{O}(l_{\rm P})$. These quantum fluctuations will lead that the geometry of spacetime vary in the local proper time scale $t_{p}$ and local proper distance scale $\lambda_{p}$. Usually, such extreme ultraviolet variations of metric have no dramatic effect on classical physics. But the appearance of horizon makes things different due to the strong redshift.
To illustrate that, we consider the quantum fluctuation happening in the near-event-horizon region $r=r_{h}+\delta r$ (measured in coordinate gauge (1)). The line element in terms of tortoise coordinate in this region then will be $\text{d}s^{2}=h(r)(-\text{d}t^{2}+\text{d}r_{*}^{2})\approx h^{\prime}(r_{h})\delta r(-\text{d}t^{2}+\text{d}r_{*}^{2})$ such that one finds $\lambda_{p}=\sqrt{h^{\prime}(r_{h})\delta r}\lambda=\sqrt{h^{\prime}(r_{h})\delta r}T$. Here $\lambda$ and $T$ denote the wavelength and period in tortoise coordinate. The picture of classical horizon would fail if one is close to it near Planck scale. Thus, the proper distance between $r=r_{h}+\delta r$ and $r=r_{h}$ can at most as same as Planck scale. This means $\delta r/\sqrt{h(r)}\approx\delta r/\sqrt{h^{\prime}(r_{h})\delta r}\sim\mathcal{O}(\lambda_{p})$. One thus finds that $T=\lambda\sim 1/h^{\prime}(r_{h})\sim\mathcal{O}(r_{h})$. Therefore, from the perspective of tortoise coordinate, near horizon local quantum fluctuations would induce metric or Regge-Wheeler potential to vary in infrared scale. It needs to emphasize again that, in the case of absent horizon or in the region far way from the horizon, the local quantum fluctuations will have no essential contributions on classical physics in infrared scale.
@@
where $W(x)$ is random function localizing around $x=0$ and $U(t)$ is a temporal random function. The function $W$ and $U$ are dimensionless functions and satisfy
@@
Figure 2: Left panel: Time evolutions of $\psi$ at $x=a$ for different values of $\epsilon$ and $a$. Right panel: Relationship between $\epsilon$ and $|a_{c}|$. The red line is $\ln|a_{c}|=-2\ln\epsilon+0.02$, which is obtained by fitting the most left 7 points.